# Evolutionary mechanisms that promote cooperation may not promote social welfare

### v1

@@ Abstract
Understanding the emergence of prosocial behaviours among self-interested individuals is an important problem in many scientific disciplines. Various mechanisms have been proposed to explain the evolution of such behaviours, primarily seeking the conditions under which a given mechanism can induce highest levels of cooperation. As these mechanisms usually involve costs that alter individual payoffs, it is however possible that aiming for highest levels of cooperation might be detrimental for social welfare – the later broadly defined as the total population payoff, taking into account all costs involved for inducing increased prosocial behaviours. Herein, by comparatively analysing the social welfare and cooperation levels obtained from stochastic evolutionary models of two well-established mechanisms of prosocial behaviour, namely, peer and institutional incentives, we demonstrate exactly that. We show that the objectives of maximising cooperation levels and the objectives of maximising social welfare are often misaligned. We argue for the need of adopting social welfare as the main optimisation objective when designing and implementing evolutionary mechanisms for social and collective goods.
@@ 1 Introduction
Since Darwin, the challenge of explaining the evolution of cooperative behavior has been actively explored across various fields, including Evolutionary Biology, Ecology, Economics, and Multi-agent Systems (Nowak, 2006b, , Sigmund,, 2010, Perc et al.,, 2017, Han, 2022a, , Paiva et al.,, 2018). Several mechanisms have been proposed to account for the evolution of cooperation, such as kin and group selection, direct and indirect reciprocity, structured populations, pre-commitments and incentives (Perc et al.,, 2017, Novak,, 2006). Therein, the emphasis is often placed on the degree or level of cooperation that a given mechanism can induce.
@@
In this paper, we demonstrate that it might be more important to optimise the social welfare, rather than focusing entirely on achieving highest levels of cooperation. Because the latter can lead to a misleading, undesirable outcome where a high cooperation level is achieved but social welfare decreases. We demonstrate these through analysing social welfare for two well-established classes of incentive mechanisms: peer and institutional incentives, for both positive (i.e. reward) and negative (i.e. punishment) types (Sigmund et al.,, 2001, Van Lange et al.,, 2014, Sasaki et al.,, 2015, Duong and Han,, 2021).
We adopt Evolutionary Game Theory (EGT) (Sigmund,, 2010, Imhof et al.,, 2005, Smith,, 1974), a well-established mathematical framework for modelling and analysing cooperative behaviours and their emergence and stability (Nowak, 2006b, , Perc et al.,, 2017). We derive close forms for the long-term expected social welfare, for population dynamics under varying mutation rates and selection intensities, which are key factors of Darwinian evolution (Nowak, 2006a, ). Our analysis is carried out using the one-shot Prisoner’s Dilemma, a well adopted game for modelling a social dilemma of cooperation (Coombs,, 1973, Sigmund,, 2010).
@@ 2.2 Evolutionary processes
where $n_{i}$ is the number of players currently adopting strategy $i$ ($i=1,\ldots,m)$. In each time step of the evolutionary process, an individual $A$ is chosen at random to update their strategy. There are two ways to do so: with probability $\mu$ (mutation rate), $A$ adopts a randomly selected strategy from the remaining $m-1$ strategies. With probability $1-\mu$, the update happens through social learning, whereby the most successful strategies tend to be imitated more often by other players (this process is equivalent to biological reproduction). That is, $A$ adopts the strategy of another, randomly chosen from the population, player $B$, with a probability given by $P_{A,B}$. A popular approach is to use the Fermi distribution from statistical physics:
$P_{A,B}=\left(1+e^{\beta(f_{A}-f_{B})}\right)^{-1},$
where $f_{A}$ ($f_{B}$) denotes the fitness of individual A (B) and $\beta$ measures the strength of the fitness contribution to the update process (a.k.a. intensity of selection). This approach leads to a unified framework for evolutionary dynamics at all intensities of selection, from random drift to imitation dynamics (Traulsen and Nowak,, 2006, Imhof et al.,, 2005).
@@
$S=|\Delta^{m}_{N}|=\binom{N+m}{m}.$
@@
$\omega_{\mathbf{n},\mathbf{n}^{\prime}}=\begin{cases}\frac{n_{i}}{S}\left(\frac{\mu}{m-1}+(1-\mu)\frac{n_{j}}{S}P_{i,j}\right)\quad\text{if}\ n^{\prime}_{i}=n_{i}-1,\ n^{\prime}_{j}=n_{j}+1,n^{\prime}_{l}=n_{l}\ \text{for }l\not\in\{i,j\},\\ 1-\sum_{j\neq i}\frac{n_{i}}{S}\left(\frac{\mu}{m-1}+(1-\mu)\frac{n_{j}}{S}P_{i,j}\right)\ \ \text{if}\ \mathbf{n}=\mathbf{n}^{\prime},\\ 0\ \ \text{otherwise.}\end{cases}$ (2)
By computing the normalised left eigenvector of the transition matrix with respect to eigenvalue $1$, we obtain the corresponding mutation-selection (stationary) distribution.
@@ 3.1 Peer incentives
We now comparatively study the efficiency of the two peer incentive approaches, for promoting cooperation and enhancing the social welfare, focusing on whether these two objectives are aligned. Indeed, Figure 1 (left column) shows that, as expected, peer punishment usually surpasses peer reward in effectively promoting cooperation, especially when selection is weaker and the impact to cost ratio of incentive is sufficiently high. However, interestingly, reward leads to higher social welfare than punishment in most cases
@@
Overall, our results have shown that, in case of peer punishment, the objective of promoting the evolution of high levels of cooperation can be detrimental for social welfare. Peer reward, on the other hand, is more efficient in promoting social welfare, even though it leads to lower levels of cooperation than punishment. The observations are robust for varying mutation rates, see Figure S3.
@@ 3.2 Institutional incentives
We now analyse the alignment between promoting cooperation and social welfare in minimal models of institutional reward and punishment (Góis et al.,, 2019, Sasaki et al.,, 2015, Duong and Han,, 2021). Namely, we consider a well-mixed population consisting of individuals playing the one-shot PD game who can adopt either C or D in the interactions.
@@
Figure 2 shows that, as expected (Duong and Han,, 2021, Han, 2022b, , Góis et al.,, 2019), both types of incentives lead to the same level of cooperation assuming the are equally effective (i.e. $a=b$) and costly. However, institutional reward leads to positive social welfare (red areas in panels a and b) for a much wider range of incentive impact and cost. For reward, even when using incentive is rather cost-inefficient, e.g. when $a=0.7$ (panel c), positive social welfare can be achieved for intermediate values of $\delta$. While for punishment, it needs to be highly efficient ($b>2$) for positive social welfare. Our observation are robust for different intensities of selection, see Supporting Information Figure S1.
@@
This is a notable observation as previous models of institutional incentives (Sasaki et al.,, 2015, Góis et al.,, 2019, Sigmund et al.,, 2010, Sasaki et al.,, 2012) do not and are unable to provide insights on how strong punishment is strong enough, focusing on promoting high levels of cooperation. In fact, previous works only consider that strong punishment is needed to ensure cooperation.
@@ 4 Discussion
However, as these mechanisms usually involve costs that alter individual payoffs, it is possible that aiming for highest levels of cooperation might be detrimental for social welfare. In this paper, using numerical simulations for stochastic evolutionary models for two important incentive mechanisms, peer and institutional incentives, we have demonstrated exactly that.
@@
Closely related to the present work are a recent body of literature on cost optimisation of institutional incentives for promoting cooperation and fairness (Han and Tran-Thanh,, 2018, Duong and Han,, 2021, Duong et al.,, 2023, Wang et al.,, 2022, Cimpeanu et al.,, 2023, Wang et al.,, 2019, Cimpeanu et al.,, 2021, Sun et al.,, 2021, Cimpeanu and Han,, 2024). These works usually consider a bi-objective optimisation problem where one aims to ensure a certain minimal level of cooperation at a smallest cost to the institution. While these works can provide insights on the optimal incentive policy, they still do not guarantee optimal social welfare. In fact, the two objectives are often misaligned (see Supporting Information, Figure S2). Moreover, we argue that focusing on the social welfare provides a more convenient, single objective optimisation problem.
@@
We analysed here two incentive mechanisms. It might be interesting to study whether maximising social welfare is aligned with maximising cooperation for other mechanisms of cooperation. For example, with kin selection, favouring related players might lead to reduction of social welfare due to unfair use of power to favour or patronage one’s relatives or friends (aka nepotism) (Wilson and Dugatkin,, 1991). For indirect reciprocity to work (Okada,, 2020), one might need to enhance transparency of reputation, e.g. via implementing institution broadcasting reputation scores (Radzvilavicius et al.,, 2021) or other costly communication mechanisms (Krellner et al.,, 2021, Santos et al.,, 2018). Thus one might need to balance between promoting cooperation and the cost of enabling it. For direct reciprocity to work (Xia et al.,, 2023), one might need to reduce noise, increase cognitive capacity (Lenaerts et al.,, 2024, Han et al.,, 2012), and improve trust (Han et al.,, 2021), which are costly and thus need to be taken into account for balancing the population social welfare. For pre-commitment to work (Nesse,, 2001), it usually requires a third party such as an institution that provides incentives (Han, 2022b, ) or broadcast commitment-based reputation (Krellner and Han,, 2023) for ensuring commitment compliance, which are costly and needed to be taken into account for enhanced social welfare.
@@
An important ongoing initiative in Artificial Intelligence (AI) research is the design and implementation of AI for social goods (Tomašev et al.,, 2020), using AI-based solutions for effectively addressing social problems. To this end, there is an emerging body of evolutionary modelling studies that address prosocial behaviours in hybrid systems of human and AI agents in co-presence (Paiva et al.,, 2018, Santos,, 2024, Han, 2022a, , Guo et al.,, 2023, Fernández Domingos et al.,, 2022, Zimmaro et al., ress, , Capraro et al.,, 2024, Shen et al.,, 2024). As developing AI is costly, it is crucial to understand what kind of AI are most conducive for prosocial behaviours, in a cost effective way. Thus, optimising the overall system payoff or social welfare is crucial for effective use of AI for social goods.

### v2

@@ Abstract
Understanding the emergence of prosocial behaviours among self-interested individuals is an important problem in many scientific disciplines. Various mechanisms have been proposed to explain the evolution of such behaviours, primarily seeking the conditions under which a given mechanism can induce highest levels of cooperation. As these mechanisms usually involve costs that alter individual payoffs, it is however possible that aiming for highest levels of cooperation might be detrimental for social welfare – the later broadly defined as the total population payoff, taking into account all costs involved for inducing increased prosocial behaviours. Herein, by comparing stochastic evolutionary models of two well-established mechanisms of prosocial behaviour—namely, peer and institutional incentives—we demonstrate that the objectives of maximising cooperation and of maximising social welfare are often misaligned. First, while peer punishment is often more effective than peer reward in promoting cooperation—especially with a higher impact-to-cost ratio—the opposite is true for social welfare. In fact, welfare typically decreases (increases) with this ratio for punishment (reward). Second, for institutional incentives, while maintaining similar levels of cooperation, rewards result in positive social welfare across a much broader range of parameters. Furthermore, both types of incentives often achieve optimal social welfare when their impact is moderate rather than maximal, indicating that careful planning is essential for costly institutional mechanisms to optimise social outcomes. These findings are consistent across varying mutation rates, selection intensities, and game configurations. Overall, we argue for the need of adopting social welfare as the main optimisation objective when designing and implementing evolutionary mechanisms for social and collective goods.
@@ 1 Introduction
Since Darwin, the challenge of explaining the evolution of cooperative behavior has been actively explored across various fields, including Evolutionary Biology, Ecology, Economics, and Multi-agent Systems (Nowak, 2006b, , Sigmund,, 2010, Perc et al.,, 2017, Han, 2022a, , Paiva et al.,, 2018, Arefin et al.,, 2020). Several mechanisms have been proposed to account for the evolution of cooperation, such as kin and group selection, direct and indirect reciprocity, structured populations, pre-commitments and incentives (Perc et al.,, 2017, Novak,, 2006). Therein, the emphasis is often placed on the degree or level of cooperation that a given mechanism can induce.
@@
In this paper, we demonstrate that it might be more important to optimise the social welfare, rather than focusing entirely on achieving highest levels of cooperation. Because the latter can lead to a misleading, undesirable outcome where a high cooperation level is achieved but social welfare decreases. We demonstrate these through analysing social welfare for two well-established classes of incentive mechanisms: peer and institutional incentives, for both positive (i.e. reward) and negative (i.e. punishment) types (Sigmund et al.,, 2001, Van Lange et al.,, 2014, Chen et al.,, 2015, Duong and Han,, 2021).
We adopt Evolutionary Game Theory (EGT) (Sigmund,, 2010, Imhof et al.,, 2005, Smith,, 1974), a well-established mathematical framework for modelling and analysing cooperative behaviours and their emergence and stability (Nowak, 2006b, , Perc et al.,, 2017). We derive close forms for the long-term expected social welfare, for population dynamics under varying mutation rates and selection intensities, which are key factors of Darwinian evolution (Nowak, 2006a, ). Our analysis is carried out using the one-shot Prisoner’s Dilemma, a well adopted game for modelling a social dilemma of cooperation (Coombs,, 1973, Sigmund,, 2010, Wang et al.,, 2015).
@@ 2.1 Prisoner’s Dilemma
The strength of the dilemma in the PD game can be varied adopting a simplified scaling approach from (Wang et al.,, 2015, Ito and Tanimoto,, 2018, Arefin et al.,, 2020). Indeed, by fixing $T-R=P-S=1$, the dilemma strength decreases when $R-P$ increases.
@@ 2.2 Evolutionary processes
where $n_{i}$ is the number of players currently adopting strategy $i$ ($i=1,\ldots,m)$. In each time step of the evolutionary process, an individual $A$ is chosen at random to update their strategy. There are two ways to do so: with probability $\mu$ (mutation rate), $A$ adopts a randomly selected strategy from the remaining $m-1$ strategies. With probability $1-\mu$, the update happens through social learning, whereby the most successful strategies tend to be imitated more often by other players (this process is equivalent to biological reproduction). That is, $A$ adopts the strategy of another, randomly chosen from the population, player $B$, with a probability given by $P_{A,B}$. A popular approach is to use the Fermi distribution from statistical physics (Szabó and Tőke,, 1998, Traulsen and Nowak,, 2006):
$P_{A,B}=\left(1+e^{\beta(f_{A}-f_{B})}\right)^{-1},$ (2)
where $f_{A}$ ($f_{B}$) denotes the fitness of individual A (B). The parameter $\beta$ represents the intensity of selection, i.e., how strongly the individuals base their decision to imitate on fitness comparison. For $\beta=0$, we obtain the limit of neutral drift – the imitation decision is random. For large $\beta$, imitation becomes increasingly deterministic. Overall, this approach leads to a unified framework for evolutionary dynamics at all intensities of selection, from random drift to imitation dynamics (Traulsen and Nowak,, 2006, Imhof et al.,, 2005).
@@
$S=|\Delta^{m}_{N}|=\binom{N+m{-1}}{m{-1}}.$ (3)
For example, in the peer incentive models we will analyse below (Section 3.1), there are three strategies, i.e. $m=3$. Thus, in that case, the Markov chain has $S=(N+2)(N+1)/2$ states. In the institutional incentive models (Section 3.2), there are two strategies ($m=2$) and the Markov chain has $S=N+1$ states.
@@
$\omega_{\mathbf{n},\mathbf{n}^{\prime}}=\begin{cases}\frac{n_{i}}{{N}}\left(\frac{\mu}{m-1}+(1-\mu)\frac{n_{j}}{{N}}P_{i,j}\right)\quad\text{if}\ n^{\prime}_{i}=n_{i}-1,\ n^{\prime}_{j}=n_{j}+1,n^{\prime}_{l}=n_{l}\ \text{for }l\not\in\{i,j\},\\ 1-\sum_{j\neq i}\frac{n_{i}}{{N}}\left(\frac{\mu}{m-1}+(1-\mu)\frac{n_{j}}{{N}}P_{i,j}\right)\ \ \text{if}\ \mathbf{n}=\mathbf{n}^{\prime},\\ 0\ \ \text{otherwise.}\end{cases}$ (4)
The first equation represents the probability that a player who chooses strategy $i$ will adopt strategy $j$. It comprises the probability when it is due to mutation, i.e. $\mu\frac{n_{i}}{{N}}\frac{1}{m-1}$ (the second and the third fractions represent the probability of choosing a strategist $i$ from the population and that of selecting strategy $j$ from the set of $m-1$ strategies other than $i$), and the probability when it is due to imitation, i.e. $(1-\mu)\frac{n_{i}}{{N}}\frac{n_{j}}{{N}}P_{i,j}$ (the second and the third fractions represent the probabilities of choosing a strategist $i$ and $j$ from the population, respectively).
Finally, by computing the normalised left eigenvector of the transition matrix with respect to eigenvalue $1$, we obtain the corresponding mutation-selection (stationary) distribution.
@@ 3.1 Peer incentives
We now comparatively study the efficiency of the two peer incentive approaches, for promoting cooperation and enhancing the social welfare, focusing on whether these two objectives are aligned. Indeed, Figure 1 (left column) shows that, as expected, peer punishment usually surpasses peer reward in effectively promoting cooperation, especially when selection is weaker and the impact to cost ratio of incentive is sufficiently high. However, interestingly, reward leads to higher social welfare than punishment in most cases. Note that when $\beta$ is very low, representing weak selection scenarios (see e.g. the last row of Figure 1, with $\beta=0.001$), the transitions between states of the Markov chain are increasingly close to random. That results in all strategies in the population having a similar frequency in stationary distribution. Thus, even when the impact to cost ratio is 0, the total level of cooperation (i.e. sum of frequencies of C and SP or SR) is close to $2/3$. On the other hand, when $\beta$ is sufficiently high, representing strong selection scenarios (see the first row of Figure 1, with $\beta=1.0$), the level of cooperation is very low even for the most efficient punishment/reward scenarios that are considered in our analysis (i.e. up to 5). The reason is that strong selection intensifies the transitions that replace a C-player with a D-player among the states in the Markov chain. Note however that if we further increase the incentive efficiency, high levels of cooperation can be achieved (See Supporting Information Figure S3).
@@
Overall, our results have shown that, in case of peer punishment, the objective of promoting the evolution of high levels of cooperation can be detrimental for social welfare. Peer reward, on the other hand, is more efficient in promoting social welfare, even though it leads to lower levels of cooperation than punishment. The observations are robust for varying mutation rates, see Supporting Information Figure S1. In fact, our observations indicate that a higher mutation rate further intensifies the inefficiency of peer punishment concerning social welfare. Moreover, we examine the robustness of the observations for varying the dilemma strength of the PD game, adopting a simplified scaling approach from (Wang et al.,, 2015, Ito and Tanimoto,, 2018, Arefin et al.,, 2020), see Supporting Information Figure S2. Again, similar observations are achieved, with a slightly improved performance of peer punishment for a weaker dilemma strength, but only when the intensity of selection is moderate (see $\beta=0.1$).
@@ 3.2 Institutional incentives
We now analyse the alignment between promoting cooperation and social welfare in minimal models of institutional reward and punishment (Góis et al.,, 2019, Chen et al.,, 2015, Duong and Han,, 2021). Namely, we consider a well-mixed population consisting of individuals playing the one-shot PD game who can adopt either C or D in the interactions.
@@
Figure 2 shows that, as expected (Duong and Han,, 2021, Han, 2022b, , Góis et al.,, 2019), both types of incentives lead to the same level of cooperation assuming they are equally effective (i.e. $a=b$) and costly. However, institutional reward leads to positive social welfare (red areas in panels a and b) for a much wider range of incentive impact and cost. For reward, even when using incentive is rather cost-inefficient, e.g. when $a=0.7$ (panel c), positive social welfare can be achieved for intermediate values of $\delta$. While for punishment, it needs to be highly efficient ($b>2$) for positive social welfare. Our observation are robust for different intensities of selection, see Supporting Information Figure S4.
@@
This is a notable observation as previous models of institutional incentives (Chen et al.,, 2015, Góis et al.,, 2019, Sigmund et al.,, 2010, Sasaki et al.,, 2012) do not and are unable to provide insights on how strong punishment is strong enough, focusing on promoting high levels of cooperation. In fact, previous works only consider that strong punishment is needed to ensure cooperation.
@@
The above-mentioned observations are also valid when considering varying dilemma strengths of the Prisoner’s Dilemma, see Supporting Information Figure S5. Notably, we also find that for weaker dilemmas (i.e. for larger $R-P$), both institutional reward and punishment result in higher levels of social welfare, while as before, reward performs better than punishment for similar effect to cost ratios (i.e. $a=b$). This further reinforces the finding that the optimal institutional incentives for cooperation and social welfare are often distinct, and that increasing the costs and impacts of incentives may be counterproductive for social welfare, regardless of whether they are rewards or punishments.
@@ 4 Discussion
However, as these mechanisms usually involve costs that alter individual payoffs, it is possible that aiming for highest levels of cooperation might be detrimental for social welfare. In this paper, using numerical simulations for stochastic evolutionary models for two important incentive mechanisms, peer and institutional incentives, we have demonstrated exactly that. For peer incentives, our finding showed that while peer punishment is often more effective than peer reward in promoting cooperation, in particular for a higher impact-to-cost ratio of the two incentives, the opposite is true for social welfare. In fact, the welfare typically decreases with this ratio for peer punishment, while it increases in case of peer reward. This is an important finding given that previous works usually focus more on peer punishment than peer reward, given the former advantage for enhancing cooperation (Sigmund et al.,, 2001, Herrmann et al.,, 2008, Sigmund et al.,, 2010, Han,, 2016, Cimpeanu and Han,, 2024, Hauert et al.,, 2007). For institutional incentives, we showed that reward fosters positive social welfare over a much wider range of parameters while maintaining similar levels of cooperation. Furthermore, both types of institutional incentives often achieve optimal social welfare when their impact is moderate rather than at the maximal level. This observation is not possible in previous models of institutional incentives as they consider only the objective of maximising cooperation (Góis et al.,, 2019, Sasaki et al.,, 2012, Sigmund,, 2010). This results indicate that careful planning is essential for costly institutional mechanisms to optimise social goods.
@@
Closely related to the present work are a recent body of literature on cost optimisation of institutional incentives for promoting cooperation and fairness (Han and Tran-Thanh,, 2018, Duong and Han,, 2021, Duong et al.,, 2023, Wang et al.,, 2022, Cimpeanu et al.,, 2023, Wang et al.,, 2019, Cimpeanu et al.,, 2021, Sun et al.,, 2021, Cimpeanu and Han,, 2024, Sun et al.,, 2023). These works usually consider a bi-objective optimisation problem where one aims to ensure a certain minimal level of cooperation at a smallest cost to the institution. While these works can provide insights on the optimal incentive policy, they still do not guarantee optimal social welfare. In fact, the two objectives are often misaligned (see Supporting Information, Figure S6). Moreover, we argue that focusing on the social welfare provides a more convenient, single objective optimisation problem.
@@
We analysed here two incentive mechanisms. It might be interesting to study whether maximising social welfare is aligned with maximising cooperation for other mechanisms of cooperation. For example, with kin selection, favouring related players might lead to reduction of social welfare due to unfair use of power to favour or patronage one’s relatives or friends (aka nepotism) (Wilson and Dugatkin,, 1991). For indirect reciprocity to work (Okada,, 2020), one might need to enhance transparency of reputation, e.g. via implementing institution broadcasting reputation scores (Radzvilavicius et al.,, 2021) or other costly communication mechanisms (Krellner et al.,, 2021, Santos et al.,, 2018). Thus one might need to balance between promoting cooperation and the cost of enabling it. For direct reciprocity to work (Xia et al.,, 2023), one might need to reduce noise, increase cognitive capacity (Lenaerts et al.,, 2024, Han et al.,, 2012), and improve trust (Han et al.,, 2021), which are costly and thus need to be taken into account for balancing the population social welfare. For pre-commitment to work (Nesse,, 2001), it usually requires a third party such as an institution that provides incentives (Han, 2022b, ) or broadcast commitment-based reputation (Krellner and Han,, 2023) for ensuring commitment compliance, which are costly and needed to be taken into account for enhanced social welfare. In the context of networked reciprocity (Nowak, 2006b, , Xia et al.,, 2023, Santos and Pacheco,, 2005), different network structures and their heterogeneity can produce varying outcomes for cooperation and social welfare. Therefore, it may be necessary to consider various costly interventions, such as state-based (Sun et al.,, 2023) or neighborhood-based incentives (Han et al.,, 2018, Cimpeanu et al.,, 2021), to achieve optimal results.
@@
We focused here on the maximisation of the population social welfare as a whole. In many domains, it can be useful to look deeper into the distribution of the social welfare across the population, as for the same total welfare, a fairer distribution of it might be more desirable. (It is noteworthy that this is not a problem in our model as in a well-mixed population, all cooperators would receive the same payoff.) For example, in the resource allocation scenarios such as those modelled by the Ultimatum and Dictator games (Hoffman et al.,, 1994, Zisis et al.,, 2015, Rand et al.,, 2013, Cimpeanu et al.,, 2021), a fair distribution between resource providers and recipients are preferred, even for a lower total welfare. In addition, a critical ongoing issue in advanced Artificial Intelligence (AI) developments is to avoid power concentration among fewer big techs (Verdegem,, 2024), in order to ensure equitable redistribution of the vast profits from advanced AI. In this domain, it would be interesting to explore mechanisms such as Windfall Clause (O’Keefe et al.,, 2020), open-source developments (Widder et al.,, 2023), regulatory markets (Clark and Hadfield,, 2019, Bova et al.,, 2024), and voluntary safety commitments (Han et al.,, 2022), as potential approaches for ensuring fair and optimal social welfare. In all these examples, one might need to balance between efficiency (i.e. the total welfare) and equity (fair distribution of the welfare). This issue has been studied extensively in operation research literature, see e.g. a survey in (Karsu and Morton,, 2015), providing suitable objective functions that combine both factors.
An important ongoing initiative AI research is the design and implementation of AI for social goods (Tomašev et al.,, 2020), using AI-based solutions for effectively addressing social problems. To this end, there is an emerging body of evolutionary modelling studies that address prosocial behaviours in hybrid systems of human and AI agents in co-presence (Paiva et al.,, 2018, Santos,, 2024, Han, 2022a, , Guo et al.,, 2023, Fernández Domingos et al.,, 2022, Zimmaro et al., ress, , Capraro et al.,, 2024, Shen et al.,, 2024). As developing AI is costly, it is crucial to understand what kind of AI are most conducive for prosocial behaviours, in a cost effective way. Thus, optimising the overall system payoff or social welfare is crucial for effective use of AI for social goods.